OpenStudy (anonymous):

determine if series converges or diverges using the comparison test: sum(n=1 to infinity)[3/(8n+7sqrt(n))]

OpenStudy (anonymous):

diverges for sure

OpenStudy (anonymous):

you can compare it to \[\sum\frac{1}{5n}\]

OpenStudy (anonymous):

since \(\sqrt{n}<n\)

OpenStudy (anonymous):

clear? or no?

OpenStudy (anonymous):

no, not really..

OpenStudy (anonymous):

ok lets take it one step at a time

OpenStudy (anonymous):

\(\sqrt{n}<n\) for any \(n\) so \(8n+7\sqrt{n}<8n+7n=15n\)

OpenStudy (anonymous):

since \(15n>8n+7\sqrt{n}\) taking reciprocals tells you \[\frac{3}{8n+7\sqrt{n}}>\frac{3}{15n}\]

OpenStudy (anonymous):

\[\sum\frac{3}{15n}=\frac{1}{5}\sum\frac{1}{n}\] the well known divergent harmonic series since \(\sum\frac{1}{n}\) diverges and since your series has larger terms, it too must diverge by the comparison test

OpenStudy (anonymous):

Thanks A Lot!!! :)

OpenStudy (anonymous):

yw